The mass of eight particles, in mg in a sample are given below.
particle No. 1 2 3 4 5 6 7 8
mass of particle 25 26 22 19 20 27 24 25
Determine the mean mass and root mean mass of the system.
1
Expert's answer
2020-09-24T11:20:46-0400

The mean mass:

"\overline{m}=\sum{\frac{m_i}{n}}=\frac{25+26+22+19+20+27+24+25}{8}=23.5"


The root mean mass:

"m_{rms}=\sqrt{\frac{25^2+26^2+22^2+19^2+20^2+27^2+24^2+25^2}{8}}" "=559.5"


Answer: 23.5; 559.5